OpenStudy (anonymous):

Pete and Lynn travel on bicycles from the same place, in opposite directions, Pete travelling 4mph faster than Lynn. After 5 hours, they are 120 miles apart. What is Lynn's rate of travel?

OpenStudy (anonymous):

what is your speed if you travel 120 miles in 5 hours?

OpenStudy (anonymous):

their combined speed = 120/5 lets say lynn travels at x mph x+x+4 = 120/5 solve for x

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

wait, how did you get that? sow the steps here

OpenStudy (anonymous):

2x+4=24 2x=20 x=10 ?

OpenStudy (anonymous):

yup, thats right. very good!

OpenStudy (anonymous):

Lol thanx

OpenStudy (anonymous):

np ::)